OpenStudy (anonymous):

three non collinear points determine a triangle how many triangles can be formed with 8 points no three of which are collinear

OpenStudy (anonymous):

This is equivalent to asking how many different combinations of 3 points you can pick from a given set of points (8 in this case), because every combination of 3 non collinear points will form a triangle. Mathematically: \( C (8,3) = \frac{8!}{5!3!} = \frac{8*7*6 }{3*2*1} = 56 \)